# DiarizationLM: Speaker Diarization Post-Processing with Large Language Models

### 07 Jan 2024 v1

@@ Abstract
In this paper, we introduce DiarizationLM, a framework to leverage large language models (LLM) to post-process the outputs from a speaker diarization system. Various goals can be achieved with the proposed framework, such as improving the readability of the diarized transcript, or reducing the word diarization error rate (WDER). In this framework, the outputs of the automatic speech recognition (ASR) and speaker diarization systems are represented as a compact textual format, which is included in the prompt to an optionally finetuned LLM. The outputs of the LLM can be used as the refined diarization results with the desired enhancement. As a post-processing step, this framework can be easily applied to any off-the-shelf ASR and speaker diarization systems without retraining existing components. Our experiments show that a finetuned PaLM 2-S model can reduce the WDER by rel. $25.9\%$ on the Fisher telephone conversation dataset, and rel. $31\%$ on the Callhome English dataset [^1] This project is currently a work in progress. We may add new experiment results to this paper soon..
@@ 3.5 LLM finetuning
To finetune the LLM, we build our training data as a collection of prompt-completion pairs. First, for each utterance, we run the ASR model and the speaker diarization system on it, and apply the orchestration module as shown in Fig. 2. This will produce the hypothesis word sequence $\mathbf{w}^{hyp}$ and hypothesis speaker sequence $\mathbf{s}^{hyp}$. From the ground truth annotations of this utterance, we build the reference word sequence $\mathbf{w}^{ref}$ and the reference speaker sequence $\mathbf{s}^{ref}$. Next, we apply the Transcript-Preserving Speaker Transfer algorithm from Section 3.4 by treating hypothesis sequences as source and reference sequences as target:
@@
Note that it is critical for the prompt and completion to use the same word sequence with different speaker sequences. This helps the LLM to focus on correcting the speaker labels without modifying the ASR transcripts.
@@ 4.1 Datasets
To finetune the LLM, we use the training subset of the Fisher corpus [40], which consists of 1,920 hours of 11,527 conversations.
@@ 4.3 Models
In our prompt builder module, we use an empty prompt prefix, and a 5-character prompt suffix " --> " (note the two spaces around the arrow). For the completions in our training data, we use a 6-character completion suffix " [eod]" (note the leading space). After processing the training data with the prompt builder module, we result in 13,430 prompt-completion pairs for training in total. The average length of a prompt is 2,371 SPM tokens, and the average length of a completion is 2,329 tokens. The LLM is trained for 1,200 steps with a batch size of 16.
@@ 4.4 Experimental results
In Table 1, we show the evaluation results of the USM + turn-to-diarize baseline together with the outputs post-processed by DiarizationLM. On the Fisher testing set, we see a rel. 25.9% improvement of WDER after post-processing with DiarizationLM. Even if we did not use any Callhome data during the LLM finetuning, on the Callhome testing set, we see a rel. 31% improvement of WDER. The WER of the USM on the two testing sets are relatively high due to domain mismatch and suboptimal annotation quality of the ground truth. However, this also demonstrated that the DiarizationLM solution provides consistent quality gains even with out-of-domain ASR and speaker diarization models.
To further demonstrate this, in Table 2, we show the results of a similar setup, but we replace the USM-based ASR model directly by the ground truth ASR transcripts from the testing sets. For these experiments, we will have WER=0%. From the table, we can still see big improvements of WDER after post-processing the diarization results by the same DiarizationLM model.
@@
Table 1: Evaluation results of the USM + turn-to-diarize baseline system and the results post-processed by DiarizationLM.
@@
WER (%) WDER (%) cpWER (%) WDER (%) cpWER (%)
Fisher 15.48 5.32 21.19 3.94 18.55
Callhome 15.36 7.72 24.39 5.33 21.47
Table 2: Evaluation results of the turn-to-diarize baseline system with reference ASR transcript (assuming WER=0%) and the results post-processed by DiarizationLM.
@@
WDER (%) cpWER (%) WDER (%) cpWER (%)
@@
Table 3: Example cases from the Fisher testing set where we see big absolute WDER reduction ($\Delta$ WDER) with DiarizationLM.
@@
Table 4: Example cases from the Callhome testing set where we see big absolute WDER reduction ($\Delta$ WDER) with DiarizationLM .
@@ 7 Conclusion
In this paper, we demonstrate that large language models (LLM) can be used to post-process speaker diarization results, achieving various goals such as improving the readability of the diarization transcript, and reducing the diarization errors. Specifically, we proposed DiarizationLM, a framework where we use a finetuned LLM to refine the results from off-the-shelf ASR and speaker diarization systems. Our experiments on Fisher and Callhome datasets show that a finetuned PaLM 2-S model can significantly reduce the word diarization error rates of typical diarization systems like turn-to-diarize.

### 16 Jan 2024 v2

@@ Abstract
In this paper, we introduce DiarizationLM, a framework to leverage large language models (LLM) to post-process the outputs from a speaker diarization system. Various goals can be achieved with the proposed framework, such as improving the readability of the diarized transcript, or reducing the word diarization error rate (WDER). In this framework, the outputs of the automatic speech recognition (ASR) and speaker diarization systems are represented as a compact textual format, which is included in the prompt to an optionally finetuned LLM. The outputs of the LLM can be used as the refined diarization results with the desired enhancement. As a post-processing step, this framework can be easily applied to any off-the-shelf ASR and speaker diarization systems without retraining existing components. Our experiments show that a finetuned PaLM 2-S model can reduce the WDER by rel. $55.5\%$ on the Fisher telephone conversation dataset, and rel. $44.9\%$ on the Callhome English dataset [^1] This project is currently a work in progress. We may add new experiment results to this paper soon..
@@ 3.5 LLM finetuning
    Avoid undesired leading or tailing text from the generated completions, such as “Here is the corrected transcript” or “We corrected the speakers for these words”;
  5. 5.
@@
To finetune the LLM, we build our training data as a collection of prompt-completion pairs. First, for each utterance, we run the ASR model and the speaker diarization system on it, and apply the orchestration module as shown in Fig. 2. This will produce the hypothesis word sequence $\mathbf{w}^{hyp}$ and hypothesis speaker sequence $\mathbf{s}^{hyp}$. From the ground truth annotations of this utterance, we build the reference word sequence $\mathbf{w}^{ref}$ and the reference speaker sequence $\mathbf{s}^{ref}$. Given these four sequences, we can build the prompts and completions in our training data with three different flavors, as introduced below.
3.5.1 Flavor 1: hyp2ora
The first flavor is named hypothesis-to-oracle, or simply hyp2ora. In this flavor, we apply the Transcript-Preserving Speaker Transfer algorithm from Section 3.4 by treating reference sequences as source and hypothesis sequences as target:
$\mathbf{s}^{ora}=\mathrm{TPST}(\mathbf{w}^{ref},\mathbf{s}^{ref},\mathbf{w}^{hyp},\mathbf{s}^{hyp}),$ (1)
where the output $\mathbf{s}^{ora}$ is the oracle hypothesis speakers transfered from the reference sequences. With $\mathbf{s}^{ora}$, the prompts and completions in our training data are created as below:
  • Prompts: The text representation of $\mathbf{w}^{hyp}$ and $\mathbf{s}^{hyp}$, with segmentation, and optionally prefix and suffix.
  • Completions: The text representation of $\mathbf{w}^{hyp}$ and $\mathbf{s}^{ora}$, with segmentation, and optionally suffix.
3.5.2 Flavor 2: deg2ref
The second flavor is named degraded-to-reference, or simply deg2ref. In this flavor, we apply the Transcript-Preserving Speaker Transfer algorithm from Section 3.4 by treating hypothesis sequences as source and reference sequences as target:
@@ 3.5.2 Flavor 2: deg2ref
3.5.3 Flavor 3: mixed
The third flavor named mixed is simply the union of the prompts and completions from the previous two flavors. When building training batches, prompt-completion pairs from the two flavors are interleaved.
Note that for all three flavors, it is critical for the prompt and completion to use the same word sequence with different speaker sequences. This helps the LLM to focus on correcting the speaker labels without modifying the ASR transcripts.
@@ 4.1 Datasets
To finetune the LLM, we use the training subset of the Fisher corpus [40], which consists of 1,920 hours of 11,527 conversations. The same train-test split of the Fisher dataset has been used in many previous works [8, 17, 35, 41]
@@ 4.2 Metrics
All three metrics reported in this paper (WER, WDER, and cpWER) are macro metrics, i.e. both numerators and denominators are aggregated on the entire dataset.
@@ 4.3 Models
In our prompt builder module, we use an empty prompt prefix, and a 5-character prompt suffix " --> " (note the two spaces around the arrow). For the completions in our training data, we use a 6-character completion suffix " [eod]" (short for “end of document”; note the leading space). After processing the training data with the prompt builder module, we result in 13,430 prompt-completion pairs for training in total. The average length of a prompt is 2,371 SPM tokens, and the average length of a completion is 2,329 tokens. The LLM is trained for 1,200 steps with a batch size of 16.
@@ 4.4 Experimental results
In Table 1, we show the evaluation results of the USM + turn-to-diarize baseline together with the outputs post-processed by DiarizationLM with three different finetuning flavors. On both datasets, we observe big improvement of both WDER and cpWER with any of the three finetuning flavors. Interesting, the biggest improvement is observed with the hyp2ora flavor, while the smallest improvement is observed with the deg2ref flavor. Specifically for hyp2ora, we see a rel. 55.5% improvement of WDER after post-processing with DiarizationLM on the Fisher testing set. Even if we did not use any Callhome data during the LLM finetuning, we see a rel. 44.9% improvement of WDER on the Callhome testing set. The WER of the USM on the two testing sets are relatively high due to domain mismatch and suboptimal annotation quality of the ground truth. However, this also demonstrated that the DiarizationLM solution provides consistent quality gains even with out-of-domain ASR and speaker diarization models.
To further demonstrate this, in Table 2, we show the results of a similar setup, but we replace the USM-based ASR model directly by the ground truth ASR transcripts from the testing sets. For these experiments, we will have WER=0%, and the hyp2ora and deg2ref flavors will be equivalent. From the table, we can still see big improvements of WDER after post-processing the diarization results by the same DiarizationLM model (i.e. deg2ref flavor in Table 1).
@@
Table 1: Evaluation results of the USM + turn-to-diarize baseline system and the results post-processed by DiarizationLM with three different finetuning flavors. All numbers are percentages.
@@
hyp2ora flavor deg2ref flavor mixed flavor
WER WDER cpWER WDER cpWER WDER cpWER WDER cpWER
Fisher 15.48 5.32 21.19 2.37 16.93 3.94 18.55 2.41 16.94
Callhome 15.36 7.72 24.39 4.25 20.22 5.33 21.47 4.76 20.84
Table 2: Evaluation results of the turn-to-diarize baseline system with reference ASR transcript (assuming WER=0%) and the results post-processed by DiarizationLM. All numbers are percentages.
@@
WDER cpWER WDER cpWER
@@
Table 3: Example cases from the Fisher testing set where we see big absolute WDER reduction ($\Delta$ WDER) with DiarizationLM (deg2ref flavor).
@@
Table 4: Example cases from the Callhome testing set where we see big absolute WDER reduction ($\Delta$ WDER) with DiarizationLM (deg2ref flavor).
@@ 7 Conclusion
In this paper, we demonstrate that large language models (LLM) can be used to post-process speaker diarization results, achieving various goals such as improving the readability of the diarization transcript, and reducing the diarization errors. Specifically, we proposed DiarizationLM, a framework where we use a finetuned LLM to refine the results from off-the-shelf ASR and speaker diarization systems. We introduced three different flavors to build the prompt-completion pairs data for finetuning the LLM. Our experiments on Fisher and Callhome datasets show that a finetuned PaLM 2-S model can drastically reduce the word diarization error rates of typical diarization systems like turn-to-diarize.